OpenStudy (anonymous):

A begins a business with Rs. 4500 and was joined afterwards by B with Rs. 5400 If the profit at the end of the year was divided in the ration 2:1, then the time of B joining was after (a) 5 months (b) 7 " (c) 8 " (d) 9 "

OpenStudy (anonymous):

Is there any formula for this? I'm new to shares.

Directrix (directrix):

(b) 7 months -------------- After reading sample problems at the link http://www.bankexam.co.in/uploads/exam_features/1008101281417227Partnershi.pdf, I decided to give this problem a shot. 4500 * 12 2 ---------- = ---- 5400 * (12-x) 1 4500 * 12 = 2 * (5400) * (12 - x) x = 7 months